A vendor bought toffees at 6 for a rupee. How many for a rupee must he sell to gain 20%

Grade:

1 Answers

Yash Baheti IIT Roorkee
9 years ago
Hi

The selling price will be = Cost price + profit = 6 + 6*(20/100) = 6 + 1.2 = 7.2

So in order to have 20% profit, he must seel the toffees at 7.2 Rs / toffee.

Or in other words : (1 / 7.2)th fraction of a toffee for a rupee.

….….….….….….….….….….….….….….….….….….….….….….….….….….…..